# A Unified Understanding and Evaluation of Steering Methods

### v1

@@ 4.3 In-depth Analysis
The evaluation protocol employed in existing work typically applies steering vectors to all examples in the test set regardless of how aligned the model would be without steering. While the steering vector intends to move negative embeddings to positive embeddings, it remains unclear how it impacts positive examples. To understand this, we empirically verify the change in performance after steering is applied only for the positive test examples, where the model assigns a higher probability to the correct answer and we report the results in Table 4. We find that the effect of steering on positive test examples is often harmful for the mean of differences and PCA-based methods. However, despite the large decreases in APC on steered positive examples for the mean of differences, the method still had the best performance in terms of both APC and ACC when considering all test examples. This suggests that the mean of differences greatly increases the APC for negative or incorrect examples, and the steering vector accomplishes the objective it was learned with, mapping negative examples to positive ones. Moreover, steering vectors can fail to generalize to positive examples of the target behavior, and more controlled applications of steering can lead to significant improvements.
Dataset CAA RepE Top PC ITI
$\Delta$APC $\Delta$APC $\Delta$APC $\Delta$APC
Coordination -9.07 -12.63 -12.92 8.75
Corrigible -$4$1.31 -3.30 10.71
One Box -6.61 1.63 1.53 12.36
Refusal -33.53 -7.59 -11.07 2.61
Self-Awareness -20.82 -19.38 -19.33 1.44
Survival -53.25 6.25 6.17 7.35
Wealth Seeking -23.98 -5.23 -5.40 9.78
Table 4: Multiple-choice behavioral evaluation only for test examples where the unsteered model assigns higher probability to the correct answer. $\Delta$APC indicates the change in APC and higher numbers are better.
When steering vectors are learned with an objective of mapping negative examples to positive examples, the training distribution of embeddings for which steering is applied is the set of negative examples. The objective does not consider the effect on positive examples, resulting in negative effects when steering is applied to such positive examples at test time. This points towards limitations of associating directions with high-level behaviors and the importance of contextualizing directions and steering vectors in terms of the data distribution. Our analysis illuminates how even in cases where the steering vector results in an overall positive effect, the steering vector actually harms performance for a large portion of the data.
@@
In Figure 3, we delve into positive steering performance using representations extracted from different layers of the LLM and from different locations within a Transformer layer. Specifically, we compare four options: (1) After the attention block, before the first residual connection; (2) After the first residual connection, before the MLP block; (3) After the MLP block, before the second residual connection; and (4) After the second residual connection, which is also known as the residual stream. We observe a notable trend that the steering performance initially increases from the top to the middle layers (e.g., 13-th layers), followed by a subsequent decline. This observation echoes prior findings that indicate representations at intermediate layers (Rimsky et al., 2024) are the most effective for steering. Furthermore, we observe that the model tends to have the best steering performance using the residual stream, which is employed for our evaluations in Section 4.1 and Section 4.2. Due to space limitation, we provide results for all datasets in Appendix C.
@@ 5 Related Works
A wide body of research addresses the problem of model alignment (Casper et al., 2023; Ji et al., 2023; Hendrycks et al., 2021; Leike et al., 2018). One of the primary approaches for model alignment is training the model on datasets consisting of human preferences on model generations (Bai et al., 2022a; Ouyang et al., 2022; Christiano et al., 2017; Ziegler et al., 2019; Yuan et al., 2023; Rafailov et al., 2023; Dong et al., 2023; Liu et al., 2023a; Song et al., 2023; Pal et al., 2024; Meng et al., 2024). A similar approach utilizes model-generated feedback to train other models for alignment (Bai et al., 2022b; Lee et al., 2023; Burns et al., 2023). There are also a wide range of works that involve detecting when a language model may generate harmful text or hallucinations (Du et al., 2024; Yi et al., 2024; Jain et al., 2023; Malinin & Gales, 2020; Kuhn et al., 2023; Duan et al., 2023; Su et al., 2024; Yin et al., 2024; Chen et al., 2024). Another line of work aims to apply modifications to the model at generation time, including steering methods (Khanov et al., 2024; Li et al., 2024; Zou et al., 2023; Rimsky et al., 2024; Liu et al., 2023b; Lee et al., 2024; Cao et al., 2024b, a; Stickland et al., 2024; Wang et al., 2024; Subramani et al., 2022; Turner et al., 2023).
Steering Model Behavior Subramani et al. (2022) introduces the idea of learning a steering vector by optimizing the vector to maximize the probability of outputting target sentences. Cao et al. (2024a) learns a steering vector by optimizing an objective that increases the ratio between the likelihood of the positive target sentence and that of a negative sentence. Other steering methods learn the steering vector by using the model’s intermediate activations on contrastive examples such as replacing “love” with “hate” (Turner et al., 2023; Li et al., 2024; Zou et al., 2023; Rimsky et al., 2024). Recent works have introduced methods that allow for more controlled applications of steering vectors by using thresholds or having multiple vectors (Lee et al., 2024; Stickland et al., 2024; Wang et al., 2024; Cao et al., 2024b; Chu et al., 2024). Tan et al. (2024) analyzes the generalization and reliability of steering vectors learned from CAA (Rimsky et al., 2024), and Singh et al. (2024) analyzes affine steering functions. Our analysis focuses on the performance and theoretical understanding behind different steering methods. In existing literature (Marks & Tegmark, 2023; Li et al., 2024; Zou et al., 2023), the comparison among different steering methods is limited to a single dataset or to simpler tasks such as classifying True/False statements. In contrast, we perform a more thorough evaluation and additionally provide theoretical reasoning to explain the performance of different methods in practice.
@@ 6 Conclusion
Our work provides the first rigorous evaluation of learning methods for steering vectors, and reveals the strengths and limitations of steering methods. Our evaluation spans multiple datasets including multiple-choice and open-ended generation tasks, explores both positive and negative steering, and verifies results across different layers. We find both theoretically and empirically that the mean difference vector results in the best performance, and that PCA-based and classifier-based methods such as those proposed in (Li et al., 2024; Zou et al., 2023) can lead to significant issues in the direction and scale of the steering vector. We provide illustrations and explanations for how and why these issues arise, fostering a deeper understanding of existing methods. These insights offer a foundation for improving the design and application of steering techniques, paving the way for more reliable and effective control of large language models.
@@
Limitations and Future Work We find through our empirical analysis that there is significant room for improvement in how steering vectors are applied with one potential avenue being more controlled and fine-grained applications of steering as seen with recent works such as (Lee et al., 2024; Stickland et al., 2024; Wang et al., 2024; Cao et al., 2024b). Additionally, while we focus on isolating the effects of steering vector design, we do not investigate how steering generalizes to new distributions at inference time. Addressing these questions requires extending the framework to account for unpredictable input distributions, which remains an open area for future research. We hope that our work illuminates the mechanisms behind steering, and provides a foundation for future development and understanding of steering model behavior.
Acknowledgement
We gratefully acknowledge Changdae Oh and Hyeong Kyu Choi for valuable comments on the draft. This work is supported by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 and IIS2331669, Office of Naval Research under grant number N00014-23-1-2643, and Philanthropic Fund from SFF.
Impact Statement
@@ Acknowledgement
Aligning language models to a set of values or human preferences is a critical research problem that addresses safety concerns around deploying machine learning models in the real world. Our research provides an in-depth evaluation and understanding of steering vectors, an approach to aligning models that are becoming widely adopted. Our findings provide practical insights for steering, showing which method has the best performance and why and the need for more controlled applications of steering. Furthermore, our analysis highlights the need for caution around associating high-level behaviors with a direction or vector. We are committed to enhancing reproducibility and broader applicability by releasing our code publicly.
@@ Appendix A Proofs
Given any joint distribution of contrastive pairs $\Pi_{+,-}$, such that the marginal distributions over positive and negative embeddings have finite means, and the objective given in (8), then the steering vector that minimizes the objective is the mean of differences:
@@
Then, we can consider $\frac{\partial\mathcal{L}}{\partial\mathbf{v}}$,
@@ A.2 Classifier Gradient Direction
Appendix B Additional Details
@@
Appendix C Additional Experiments
C.1 GPT-4 Raw Scores
Dataset CAA RepE Top PC ITI
Score Score Score Score
Coordination 0.5 0.24 0.16 0.36
Corrigible 7.04 5.16 5.66 7.82
One Box 8.24 7.50 7.60 7.39
Refusal 10 4.54 4.42 10
Self-Awareness 9.16 9.00 9.00 9.00
Survival 5.30 5.40 5.34 5.98
Wealth Seeking 8.90 8.94 8.82 9.38
Table 5: Open-ended behavioral evaluation. For all methods, steering is performed at layer 13. The scores are ratings given by GPT-4 based on how well the response aligns with the behavior being steered. For positive steering, higher numbers are better.
Dataset CAA RepE Top PC ITI
Score Score Score Score
Coordination 0.28 0.20 0.08 0.46
Corrigible 4.76 5.32 5.16 0.24
One Box 5.58 7.96 6.67 7.30
Refusal 9.6 4.58 4.34 10
Self-Awareness 8.40 9.00 9.00 9.00
Survival 4.41 5.26 5.24 4.52
Wealth Seeking 8.46 8.82 8.82 6.86
Table 6: Open-ended behavioral evaluation. For all methods, steering is performed at layer 13. The scores are ratings given by GPT-4 based on how well the response aligns with the behavior being steered. For negative steering, lower numbers are better.
C.2 Performance Across Embedding Location

### v2

@@ 4.3 In-depth Analysis
The evaluation protocol employed in existing work typically applies steering vectors to all examples in the test set regardless of how aligned the model would be without steering. While the steering vector intends to move negative embeddings to positive embeddings, it remains unclear how it impacts positive examples. To understand this, we empirically verify the change in performance after steering is applied only for the positive test examples, where the model assigns a higher probability to the correct answer and we report the results in Table 6. We find that the effect of steering on positive test examples is often harmful for the PCA-based methods, and the effect of steering on positive test examples is smaller than the average change in APC over all examples.
Dataset MoD PoD PoE CoE
$\DeltaAPC Δ\DeltaAPC$ $\DeltaAPC Δ\DeltaAPC$
Coordination 9.04 -12.63 -12.92 8.75
Corrigible 9.81 -1.31 -3.30 10.71
One Box 12.27 1.63 1.53 12.36
Refusal 1.29 -7.59 -11.07 2.61
Self-Awareness -6.80 -19.38 -19.33 1.44
Survival 3.08 6.25 6.17 7.35
Wealth Seeking 9.55 -5.23 -5.40 9.78
[…]
When steering vectors are learned with an objective of mapping negative examples to positive examples, the training distribution of embeddings for which steering is applied is the set of negative examples. The objective does not consider the effect on positive examples, potentially resulting in negative effects when steering is applied to such positive examples at test time. This points towards limitations of associating directions with high-level behaviors and the importance of contextualizing directions and steering vectors in terms of the data distribution. Our analysis illuminates how even in cases where the steering vector results in an overall positive effect, the steering vector can actually harm performance for a large portion of the data.
@@
In Figure 3, we delve into positive steering performance using representations extracted from different layers of the LLM and from different locations within a Transformer layer. Specifically, we compare four options: (1) After the attention block, before the first residual connection; (2) After the first residual connection, before the MLP block; (3) After the MLP block, before the second residual connection; and (4) After the second residual connection, which is also known as the residual stream. We observe a notable trend that the steering performance initially increases from the top to the middle layers (e.g., 13-th layers), followed by a subsequent decline. This observation echoes prior findings that indicate representations at intermediate layers [41] are the most effective for steering. Furthermore, we observe that the model tends to have the best steering performance using the residual stream, which is employed for our evaluations in Section 4.1 and Section 4.2. We provide further analysis on Mistral-7B-Instruct-v0.3 [22] and Llama-3.1-8B-Instruct [17] in Appendix B.
@@ 5 Related Works
A wide body of research addresses the problem of model alignment [9, 21, 19, 27]. One of the primary approaches for model alignment is training the model on datasets consisting of human preferences on model generations [2, 36, 12, 56, 55, 40, 14, 29, 44, 37, 34]. A similar approach utilizes model-generated feedback to train other models for alignment [3, 26, 6]. There are also a wide range of works that involve detecting when a language model may generate harmful text or hallucinations [15, 53, 20, 32, 24, 16, 46, 54, 10]. Another line of work aims to apply modifications to the model at generation time, including steering methods [23, 28, 57, 41, 30, 25, 8, 7, 45, 51, 47, 50].
Steering Model Behavior Subramani et al. [47] introduces the idea of learning a steering vector by optimizing the vector to maximize the probability of outputting target sentences. Cao et al. [7] learns a steering vector by optimizing an objective that increases the ratio between the likelihood of the positive target sentence and that of a negative sentence. [52, 31] learn to steer based directly on the output generation, and [18] learns to align the neurons with variables from some known causal model. Other steering methods learn the steering vector by using the model’s intermediate activations on contrastive examples such as replacing “love” with “hate” [50, 28, 57, 41]. Recent works have introduced methods that allow for more controlled applications of steering vectors by using thresholds or having multiple vectors [25, 45, 51, 8, 13, 11]. Other recent works have consider more general linear steering methods [42, 11, 4]. Tan et al. [48] analyzes the generalization and reliability of steering vectors learned from CAA [41], and Pres et al. [39] presents guidelines and a methodology for more reliable evaluation of steering methods. Singh et al. [43] analyzes affine steering functions, and Belrose [5] presents an analysis on worst-case behavior. While these show how mean of differences can be optimal, Singh et al. [43] considers an objective of minimizing the change in embeddings while Belrose et al. [4] focus on a worst-case objective. Our analysis focuses on the average sample-wise performance and theoretical understanding behind different steering methods. In existing literature [33, 28, 57], the comparison among different steering methods is limited to a single dataset or to simpler tasks such as classifying True/False statements. In contrast, we perform a more thorough evaluation and additionally provide theoretical reasoning to explain the performance of different methods in practice.
@@ 6 Conclusion
Our work provides the first rigorous evaluation of learning methods for steering vectors, and reveals the strengths and limitations of steering methods. Our evaluation spans multiple datasets including multiple-choice and open-ended generation tasks, explores both positive and negative steering, and verifies results across different layers. We find both theoretically and empirically that the mean difference vector results in the best performance, and that PCA-based and classifier-based methods such as those proposed in [28, 57] can lead to significant issues in the direction and scale of the steering vector. We provide illustrations and explanations for how and why these issues arise, fostering a deeper understanding of existing methods. These insights offer a foundation for improving the design and application of steering techniques, paving the way for more reliable and effective control of large language models.
@@
Limitations and Future Work We find through our empirical analysis that there is significant room for improvement in how steering vectors are applied with one potential avenue being more controlled and fine-grained applications of steering as seen with recent works such as [25, 45, 51, 8, 11, 4, 43, 42]. Additionally, while we focus on isolating the effects of steering vector design, we do not investigate how steering generalizes to new distributions at inference time. Addressing these questions requires extending the framework to account for unpredictable input distributions, which remains an open area for future research. We hope that our work illuminates the mechanisms behind steering, and provides a foundation for future development and understanding of steering model behavior.
Broader Impact Statement
@@
Aligning language models to a set of values or human preferences is a critical research problem that addresses safety concerns around deploying machine learning models in the real world. Our research provides an in-depth evaluation and understanding of steering vectors, an approach to aligning models that are becoming widely adopted. Our findings provide practical insights for steering, showing which method has the best performance and why and the need for more controlled applications of steering. Furthermore, our analysis highlights the need for caution around associating high-level behaviors with a direction or vector.
Acknowledgments
We thank Changdae Oh and Hyeong Kyu Choi for their valuable comments on the manuscript. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Shawn Im is also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2137424. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Support was also provided by the Graduate School and the Office of the Vice Chancellor for Research at the University of Wisconsin-Madison with funding from the Wisconsin Alumni Research Foundation.
@@ Appendix A Proofs
Given any joint distribution of contrastive pairs $\Pi_{+,-}$, such that the marginal distributions over positive and negative embeddings have finite means, and the objective given in equation 8, then the steering vector that minimizes the objective is the mean of differences:
@@
Then, we can consider $\frac{\partial\mathcal{L}}{\partial\mathbf{v}}$, and as we have finite means, we have
@@ A.2 Classifier Gradient Direction
Appendix B Additional Results
Evaluation on other models.
We provide multiple-choice behavior evaluations for different models including Mistral-7B-Instruct-v0.3 [22] and Llama-3.1-8B-Instruct [17] which can be seen in Tables 7 and 8 respectively. We also explore steering performance using different layers and locations with a layer for each of these models in Figures 4 and 5.
Dataset MoD PoD PoE CoE
Coordination 52.70 54 35.73 38 29.88 18 6.78 6
Corrigible 89.34 92 60.86 62 72.53 72 62.27 64
One Box Tendency 61.43 62 46.91 46 47.30 48 44.19 40
Refusal 95.84 96 84.75 84 90.47 90 86.86 88
Self-Awareness 76.52 76 58.68 60 43.84 42 39.24 36
Survival Instincts 55.02 54 54.91 58 49.28 52 40.52 40
Wealth Seeking 90.28 92 73.20 74 74.85 76 73.54 74
Table 7: Multiple-choice behavioral evaluation for Mistral-7B-Instruct-v0.3. For all methods, steering is performed at layer 13. APC scores are average token probabilities of the correct answer letter. ACC indicates the average accuracy. For positive steering, higher numbers are better.
Dataset MoD PoD PoE CoE
Coordination 58.01 68 52.51 56 52.48 54 51.80 60
Corrigible 57.27 50 48.01 48 48.08 48 53.03 60
One Box Tendency 56.39 52 52.31 58 52.43 58 50.66 52
Refusal 67.22 80 51.87 58 54.00 60 56.99 68
Self-Awareness 56.06 60 48.51 48 48.93 48 44.02 32
Survival Instincts 56.12 58 50.39 56 50.94 58 46.33 54
Wealth Seeking 66.43 80 52.08 62 52.32 62 53.79 60
Table 8: Multiple-choice behavioral evaluation for Llama-3.1-8B-Instruct. For all methods, steering is performed at layer 16. APC scores are average token probabilities of the correct answer letter. ACC indicates the average accuracy. For positive steering, higher numbers are better.
Figure 4: Performance of Mean of Differences using different layers and locations for Mistral-7B-Instruct-v0.3. We report the APC in the $y$-axis.
Figure 5: Performance of Mean of Differences using different layers and locations for Llama-3.1-8B-Instruct. We report the APC in the $y$-axis.
Appendix C Additional Details